# SkyScenes: A Synthetic Dataset for Aerial Scene Understanding

### v1

@@ Abstract
Real-world aerial scene understanding is limited by a lack of datasets that contain densely annotated images curated under a diverse set of conditions. Due to inherent challenges in obtaining such images in controlled real-world settings, we present SkyScenes, a synthetic dataset of densely annotated aerial images captured from Unmanned Aerial Vehicle (UAV) perspectives. We carefully curate SkyScenes images from Carla to comprehensively capture diversity across layout (urban and rural maps), weather conditions, times of day, pitch angles and altitudes with corresponding semantic, instance and depth annotations. Through our experiments using SkyScenes, we show that (1) Models trained on SkyScenes generalize well to different real-world scenarios, (2) augmenting training on real images with SkyScenes data can improve real-world performance, (3) controlled variations in SkyScenes can offer insights into how models respond to changes in viewpoint conditions, and (4) incorporating additional sensor modalities (depth) can improve aerial scene understanding.
| Dataset | Reproducibility | Diversity | Annotations | Image Capture | Scale | |||||||
| Metadata | Contr. Var. | Town | Daytime | Weather | Semantic | Instance | Depth | Altitude | Perspective | Resolution | ||
@@
| 1 UAVid [19] | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | Med | Obl. | K | |
| 2 AeroScapes [22] | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ | (Low, Med) | (Obl., Nad.) | K | |
| 3 ICG Drone [14] | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ | Low | Nad. | K | |
@@
| 4 MidAir [9] | ✓ | partial | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | Low | (Obl., Nad.) | K | |
| 5 VALID [2] | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | (Low, Med, High) | Nad. | K | |
| 6 Espada [17] | ✗ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | (Med, High) | Nad. | K | |
| 7 TartanAir [35] | ✓ | * | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | Low | (Fwd., Obl.) | M | |
| 8 UrbanScene3D [16] | * | * | ✓ | * | * | ✗ | ✗ | ✗ | Med | Obl. | K | |
| 9 SynthAer [30] | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | (Low, Med) | Obl. | K | |
| 10 SynDrone [25] | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | (Low, Med, High) | (Obl., Nad.) | K | |
| 11 SkyScenes | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | (Low, Med, High) | (Fwd., Obl., Nad.) | K | |
1 Introduction
We introduce SkyScenes, a densely annotated dataset of synthetic scenes captured from aerial (UAV) viewpoints under diverse layout (urban and rural), weather, daytime, pitch and altitude conditions. The setting of outdoor aerial imagery provides unique challenges for scene understanding: variability in altitude and angle of image capture, skewed representation for classes with smaller object sizes (humans, vehicles), size and occlusion variations of object classes in the same image, changes in weather or daytime conditions, etc. Naturally, training effective aerial scene-understanding models requires access to large-scale annotated exemplar data that have been carefully curated under diverse conditions. Capturing such images not only allows training models that can be robust to anticipated test-time variations but also allows assessing model susceptibility to changing conditions. However, carefully curating and annotating such images in the real-world can be prohibitively expensive due to various reasons. First, densely annotating high-resolution aerial images is expensive – for instance, densely annotating a single K image in UAVid [19] can take up to hours! Second, any effort to expand the real set to include widespread variations (weather, time of day, pitch, altitude) would be uncontrolled (i.e., can’t guarantee the same viewpoint under different conditions as real world is not static) and additionally would require re-annotating newly captured frames. Synthetic data curated from simulators can help counter both of these issues as – (1) labels are automatic and cheap to obtain and (2) it is possible to recreate the same viewpoint (with the same actor instances – vehicles, humans, etc. in the scene) under differing conditions.
Unlike synthetic ground plane view datasets (especially for autonomous driving [29, 5, 37, 21, 26]), synthetic datasets for aerial imagery (see Table. 1, rows -) relatively have received less attention [9, 2, 17, 35, 16, 30, 25] Existing synthetic datasets for aerial imagery can be lacking in a few different aspects – complementary metadata to reproduce existing frame viewpoints under different conditions, limited diversity, availability of dense annotations for a wide vocabulary of classes and image capture (height, pitch) conditions (see Table. 1 for an exhaustive summary). We cover all these aspects by introducing SkyScenes, a synthetic dataset containing K densely annotated aerial scenes captured from Carla [7]. We curate SkyScenes images by re-purposing the Carla [7] simulator for aerial viewpoints by re-positioning the agent camera to a desired altitude and pitch to obtain an oblique perspective. SkyScenes images are coupled with dense semantic, instance segmentation ( classes) and depth annotations.
We carefully curate SkyScenes images by procedurally teleoperating the aerially situated camera through distinct map layouts across different weather and daytime conditions each over a combination of altitude and pitch variations (see Fig. 1 for examples). While doing so, we keep several important desiderata in mind. First, we ensure that stored snapshots are not correlated, to promote diverse viewpoints within a town and facilitate model training. Second, we store all metadata associated with the position of actors, camera, and other scene elements to be able to reproduce the same viewpoints under different weather and daytime conditions. Thirdly, we ensure that the generated data is physically realistic. This involves introducing variations in sensor locations, such as adding jitter to specified height and pitch values, to mimic real-world conditions.111Moreover, through rigorous validations, we ensure this process is consistent and yields error-free re-generations. See Sec. 3.1 Finally, since Carla [7] by default does not spawn a lot of pedestrians in a scene, we propose an algorithm to ensure adequate representation of humans in the scene while curating images (see Sec. 3.1 and Sec. 3.2 for a detailed discussion).
Empirically, we demonstrate the utility of SkyScenes in several different ways. First, we show that SkyScenes is a “good” pre-training dataset for real-world aerial scene understanding by – (1) demonstrating that models trained on SkyScenes generalize well to multiple real-world datasets and (2) that SkyScenes pretraining can help reduce data requirements from the real-world (improved real-world performance in low-shot regimes). Second, we show that controlled variations in SkyScenes can serve as a diagnostic test-bed to assess model sensitivity to weather, daytime, pitch, altitude, and layout conditions – by testing SkyScenes trained models in unseen SkyScenes conditions. Finally, we show that SkyScenes can enable developing multi-modal segmentation models with improved aerial-scene understanding capabilities when additional sensors, such as Depth, are available. To summarize, we make the following contributions:
We introduce, SkyScenes, a densely-annotated dataset of synthetic aerial images. SkyScenes contains images from different altitude and pitch settings, encompassing different layouts, weather, and daytime conditions with corresponding dense annotations and viewpoint metadata.
We demonstrate that SkyScenes pre-trained models generalize well to real-world scenes and that SkyScenes data can effectively augment real-world training data for improved performance.
We show that SkyScenes alone can serve as a diagnostic test-bed to assess model sensitivity to changing weather, daytime, pitch, altitude and layout conditions.
Finally, we show that incorporating additional modalities (depth) while training aerial scene-understanding models can improve aerial scene recognition, enabling development of multi-modal segmentation models.
2 Related Work
@@
Ground-view Synthetic Datasets. Real world ground-view scene-understanding datasets (Cityscapes [5], Mapillary [21], BDD-100K [37], Dark Zurich [28]) fail to capture the full range of variations that exist in the world. Synthetic data is a popular alternative for generating diverse and bountiful views. GTAV [29], Synthia [26] and VisDA-C [23] are some of the widely-used synthetic datasets. These datasets can be curated using underlying simulators, such as the GTAV [29] game engine and Carla [7] simulator, and offer a cost-effective and scalable way to generate large amounts of labeled data under diverse conditions. Similar to SELMA [33] and SHIFT [32], we use Carla [7] as the underlying simulator for SkyScenes.
@@
Real-World Aerial Datasets. To support remote sensing applications, it is crucial to have access to datasets that offer aerial-specific views. Datasets such as GID [34], DeepGlobe [6], ISPRS2D [27], and FloodNet [24] primarily provide nadir perspectives and are designed for scene-recognition and understanding tasks. However, this study specifically focuses on lower altitudes, which are more relevant to UAVs, enabling object identification. Unfortunately, there is a scarcity of high-resolution real-world datasets based on UAV imagery that emphasize on object identification. Many existing urban scene datasets, like Aeroscapes [22], UAVid [19], VDD [1], UDD [4], UAVDT [8], VisDrone [38], Semantic Drones [14] and others, suffer from limited sizes and a lack of diverse images under different conditions. This limitation raises concerns regarding model robustness and generalization.
@@
Synthetic Aerial Datasets. Simulators can facilitate affordable, reliable, and quick collection of large synthetic aerial datasets, which aids in fast prototyping, improves real-world performance by enhancing robustness, and enables controlled studies on varied conditions. One such high-fidelity simulator, AirSim [31], used for the development and testing of autonomous systems (in particular, aerial vehicles), is the foundation of several synthetic UAV-based datasets like MidAir [9], Espada [17], Tartan Air [35], UrbanScene3d [16] and VALID [2]. Carla [7] is another such open-source simulator that is the foundation of datasets like SynDrone[25]. However, these datasets fall short in capturing real-world irregularities, lack deterministic re-generation capabilities, controlled diversity in weather and daytime conditions, and exhibit skewed representation for certain classes (differences summarized in Table. 1). This restricts their ability to generalize well to real-world datasets and their usage as a diagnostic tool for studying the controlled effect of diversity on the performance of computer vision perception tasks. To enable such studies, SkyScenes offers images featuring varied scenes, diverse weather, daytime, altitude, and pitch variations while incorporating real-world irregularities and addressing skewed class representation along with simultaneous depth, semantic, and instance segmentation annotations.
@@ 3 SkyScenes
We curate SkyScenes using, Carla [7]222https://carla.org/ , which is a flexible and realistic open-source autonomous vehicle simulator. The simulator offers a wide range of sensors, environmental configurations, and varying rendering configurations.
As noted earlier, we take several important considerations into account while curating SkyScenes images. These include strategies for obtaining diverse synthetic data and embedding real-world irregularities, avoiding correlated images, addressing skewed class representations and more. In this section, we first discuss such desiderata, and then describe our procedural image curation algorithm. Finally, we describe different aspects of the curated dataset.
@@
Before delving into the image curation pipeline, we first outline a set of desiderata taken into account while curating synthetic aerial images in SkyScenes.
@@
Adequate Height Variations: Aerial images are captured at different altitudes to meet specific needs. Lower altitudes (- m) are optimal for high-resolution photography and detailed inspections. Altitudes ranging from m-m strike a balance between fine-grained detail and a broader perspective, making them ideal for surveillance. Altitudes above m are suitable for capturing extensive areas, making them ideal for surveying and mapping. Existing datasets (synthetic or real) often focus on “specific” altitude ranges (see Table. 1, Image Capture columns), limiting their adaptability to different scenarios. With SkyScenes, our aim is to provide flexibility in altitude sampling, thus accommodating various real-world requirements. We curate SkyScenes images at heights of m, m and m. Additionally, recognizing imperfections in real-world actuation, we induce slight jitter in the height values (jitter ) to simulate realistic data sampling.
2.
Adequate Pitch Variations: Similar to height, aerial images can be captured from primary perspectives or pitch angles (): nadir (), oblique (), or forward () views (see Table. 1, Image Capture columns). The nadir view (directly perpendicular to the ground plane), preserves object scale while forward views are well-suited for tasks like UAV navigation and obstacle detection. Oblique views, on the other hand, capture objects from a side profile, aiding object recognition and providing valuable context and depth perspective often lost in nadir and forward views. To ensure widespread utility, SkyScenes data generation process is designed to support all these viewing angles, with a particular emphasis on oblique views (most common one). Similar to height, pitch variations allow models trained on SkyScenes to generalize to different real-world conditions. We use for oblique-views and introduce jitter (jitter ) to mimic real-world data sampling.
@@
4.
Adequate Weather & Daytime Variations: Training robust perception models using SkyScenes that generalize to unforeseen environmental conditions necessitates the curation of annotated images encompassing various weather and daytime scenarios. To accomplish this, we generate SkyScenes images from identical viewpoints under different variations – ClearNoon, ClearSunset, MidRainNoon, ClearNight and CloudyNoon.333Note that Carla [7] provides such conditions but we use only such conditions in this preliminary version of SkyScenes. Generating images in different conditions from the same perspectives allows us to (1) leverage diverse data for improved generalization and (2) systematically investigate the susceptibility of trained models to variations in daytime and weather conditions.
5.
@@
Figure 2: Ground-View vs Aerial-View Pixel Proportions. For a subset of commonly annotated classes across Cityscapes [5] (red), UAVid [19] (dark blue) and SkyScenes (light blue), we show the percentage of pixels occupied by different classes. Aerial scenes (in UAVid) have significant under-representation of tail classes (vehicle, human). SkyScenes image curation (for the same viewpoint settings as UAVid) helps counter this discrepancy via synthetic aerial scenes. -
6.
Viewpoint Reproducibility: Critical to understanding how models respond to changing conditions is the ability to evaluate them under scenarios where only one variable is altered. However, any effort to do so in the real-world would be uncontrolled, due to it’s dynamic (constantly changing) nature. In contrast, simulated data allows us to do so by providing control over image generation conditions. Unlike certain existing aerial datasets that do not support this feature (see Table. 1), we do so in SkyScenes by additionally storing comprehensive metadata for each viewpoint (and image), including details about camera world coordinates, orientation and all movable / immovable actors and objects in the scene. We couple this with rigorous consistency checks for image generation that verify the number of actors, their location, sensor height, pitch, etc. This meticulous approach enables us to reproduce the same viewpoint under multiple conditions effortlessly.
7 .
Adequate Representation of Tail Classes: Unlike ground-view datasets, pixel distribution of classes in aerial images is substantially more long-tailed (see Fig. 2; classes with smaller object size, humans), making visual recognition tasks harder. This problem is further exacerbated by differing class distributions across synthetic and real data (see Fig. 5) – especially severe for instances of human class. To counter this, we consider structured spawning of humans to ensure adequate representation.

### v2

@@ Abstract
Real-world aerial scene understanding is limited by a lack of datasets that contain densely annotated images curated under a diverse set of conditions. Due to inherent challenges in obtaining such images in controlled real-world settings, we present SkyScenes, a synthetic dataset of densely annotated aerial images captured from Unmanned Aerial Vehicle (UAV) perspectives. We carefully curate SkyScenes images from Carla to comprehensively capture diversity across layouts (urban and rural maps), weather conditions, times of day, pitch angles and altitudes with corresponding semantic, instance and depth annotations. Through our experiments using SkyScenes, we show that (1) models trained on SkyScenes generalize well to different real-world scenarios, (2) augmenting training on real images with SkyScenes data can improve real-world performance, (3) controlled variations in SkyScenes can offer insights into how models respond to changes in viewpoint conditions (height and pitch), weather and time of day, and (4) incorporating additional sensor modalities (depth) can improve aerial scene understanding. Our dataset and associated generation code are publicly available at: https://hoffman-group.github.io/SkyScenes/
Keywords:
Aerial scene understanding, Synthetic-to-Real generalization, Segmentation, Domain Generalization, Synthetic Data
1 Introduction
Aerial imagery provides a unique perspective that is invaluable for a wide range of applications, including surveillance [33, 39], mapping [37, 3], urban planning [24, 13], environmental monitoring [7, 29], and disaster response [31, 21]. These applications rely on accurate and detailed analysis of aerial images to make informed decisions and effectively address various challenges. Naturally, training effective aerial scene-understanding models requires access to large-scale annotated exemplar data that have been carefully curated under diverse conditions. Capturing such images not only allows training models that can be robust to anticipated test-time variations but also allows assessing model susceptibility to changing conditions.
However, carefully curating and annotating such images in the real-world can be prohibitively expensive due to various reasons. First, densely annotating every pixel of high-resolution real-world aerial images is expensive – for instance, densely annotating a single K image in UAVid [27] can take up to hours! Second, although diversity in training data is vital for developing robust generalization algorithms and for model sensitivity assessment, expanding the real set to include widespread variations (weather, time of day, pitch, altitude) would be uncontrolled (i.e., we can’t guarantee the same viewpoint under different conditions as the real world is not static), and hence would require re-annotating newly captured frames. Synthetic data curated from simulators can help counter both these issues as (1) labels are automatic and cheap to obtain and (2) it is possible to recreate same viewpoint (with same scene layout and actor instances, vehicles, humans, etc. in the scene) under differing conditions.
The unique challenges introduced by outdoor aerial imagery setting such as variability in altitude and angle of image capture, skewed representation for classes with smaller object sizes (humans, vehicles), size and occlusion variations of object classes in the same image, etc., make it a relatively difficult problem compared to ground-view imagery setting. This is illustrated by the observation that the current developments in syn-to-real generalization methods [54, 16] have resulted in a significant reduction in the syn-to-real generalization gap in ground-view settings to as minimal as 20%111metric of choice: mIoU., compared to the comparatively higher margin observed in aerial imagery scenarios, reaching up to 50% (see Sec. 0.D.3, 0.D.4, 0.D.5 in appendix). Unlike synthetic ground plane view datasets (especially for autonomous driving [46, 8, 56, 32, 43]), synthetic datasets for aerial imagery (see Table. 1, rows -) have received relatively less attention [12, 5, 25, 53, 23, 47, 42]. While existing synthetic aerial imagery datasets have tried to close this gap, they are often found lacking in a few different aspects – complementary metadata to reproduce existing frame viewpoints under different conditions, limited diversity, availability of dense annotations for a wide vocabulary of classes and image capture (height, pitch) conditions (see Table. 1 for an exhaustive summary). These synthetic aerial datasets rarely allow for reproducing the exact viewpoint with different variations based on detailed scene metadata (see Table.1 Controlled Variations column), a key aspect for evaluating deep learning models’ responses to changing conditions and assessing sensitivity to single-variable alterations (e.g., weather, time of day, sensor angle).
We cover all these aspects by introducing SkyScenes, a synthetic dataset containing k densely annotated aerial scenes, leveraging the Carla [10] simulator to capture diverse layout (urban and rural), weather, daytime, pitch and altitude conditions. Our pipeline meticulously generates scenes with precise actor locations and orientations, ensuring each scene’s reproducibility with a comprehensive metadata store and rigorous consistency checks.
SkyScenes encompasses detailed semantic, instance segmentation ( classes), and depth annotations across distinct map layouts across different weather and daytime conditions each over a combination of altitude and pitch variations (see Fig. 1 for examples). While doing so, we keep several important desiderata in mind. First, we ensure that stored snapshots are not correlated, to promote diverse viewpoints within a town and facilitate model training. Second, we store all metadata associated with the position of actors, camera, and other scene elements to be able to reproduce the same viewpoints under different weather and daytime conditions. Thirdly, we ensure that the generated data mimics real-world imperfections by introducing variations in sensor locations, such as adding jitter to specified height and pitch values.222Moreover, through rigorous validations, we ensure this process is consistent and yields error-free re-generations. See Sec. 3.1. Finally, since Carla [10] by default does not spawn a lot of pedestrians in a scene, we propose an algorithm to ensure adequate representation of humans in the scene while curating images (see Sec. 3.1 and Sec. 3.2 for a detailed discussion).
Our experiments across different real datasets and increasingly competitive semantic segmentation architectures consistently demonstrate that SkyScenes outperforms its closest synthetic counterpart dataset, SynDrone [42]. However, despite these dataset-level improvements, the syn-to-real generalization gap persists, indicating that algorithmic improvements developed for ground-view imagery fail to translate effectively to aerial imagery. This underscores the urgent need for specialized algorithmic development in this area.
Empirically, we demonstrate the utility of SkyScenes in several different ways. First, we show that SkyScenes is a valuable pre-training dataset for real-world aerial scene understanding by, (1) demonstrating that models trained on SkyScenes generalize well to multiple real-world datasets and (2) demonstrating that SkyScenes pretraining improves real-world performance in low-shot regimes. Second, we show that controlled variations in SkyScenes can serve as a diagnostic test-bed to assess model sensitivity to weather, daytime, pitch, altitude, and layout conditions – by testing SkyScenes trained models in unseen SkyScenes conditions. Finally, we show that SkyScenes can enable developing multi-modal segmentation models with improved aerial-scene understanding capabilities when additional sensors, such as Depth, are available. To summarize, we make the following contributions:
We introduce, SkyScenes, a densely-annotated dataset of k synthetic aerial images. SkyScenes contains images from different altitude and pitch settings, encompassing different layouts, weather, and daytime conditions with corresponding dense annotations and viewpoint metadata.
We demonstrate that SkyScenes pre-trained models generalize well to real-world scenes and that SkyScenes data can effectively augment real-world training data for improved performance. We also bring attention to the point that while the synthetic-to-real gap has considerably narrowed for ground-view datasets, the same algorithms are unable to bridge this gap in aerial imagery.
We show that our unique ability to generate controlled variations enables SkyScenes to serve as a diagnostic test-bed to assess model sensitivity to changing weather, daytime, pitch, altitude, and layout conditions.
Finally, we show that incorporating additional modalities (depth) while training aerial scene-understanding models can improve aerial scene recognition, enabling further development of multi-modal segmentation models.
2 Related Work
| Dataset | Controlled Variations | Diversity | Annotation Diversity | Altitude | Perspective | Resolution | Scale | |||||
| Town | Daytime | Weather | ||||||||||
@@ 2 Related Work
| 1 UAVid [27] | ✗ | ✗ | ✗ | ✗ | S | Med | Obl. | k | ||||
| 2 AeroScapes [34] | ✗ | ✗ | ✗ | ✗ | S | (Low, Med) | (Obl., Nad.) | k | ||||
| 3 ICG Drone [20] | ✗ | ✗ | ✗ | ✗ | S | Low | Nad. | k | ||||
@@
| 4 Espada [25] | ✗ | ✓ | ✗ | ✗ | D | (Med, High) | Nad. | k | ||||
| 5 UrbanScene3D [23] | ✗ | ✓ | ✗ | ✗ | - | Med | Obl. | k | ||||
| 6 SynthAer [47] | ✓ | ✗ | ✓ | ✗ | S | (Low, Med) | Obl. | k | ||||
| 7 MidAir [12] | ✗ | ✓ | ✓ | ✓ | (S,D) | Low | (Obl., Nad.) | k | ||||
| 8 TartanAir [53] | ✗ | ✓ | ✓ | ✓ | (S,D) | Low | (Fwd., Obl.) | M | ||||
| 9 VALID [5] | ✗ | ✓ | ✓ | ✓ | (S,I,D) | (Low, Med, High) | Nad. | k | ||||
| 10 SynDrone [42] | ✗ | ✓ | ✗ | ✗ | (S,D) | (Low, Med, High) | (Obl., Nad.) | k | ||||
| 11 SkyScenes | ✓ | ✓ | ✓ | ✓ | (S,I,D) | (Low, Med, High) | (Fwd., Obl., Nad.) | k | ||||
@@
Ground-view Synthetic Datasets. Real-world ground-view scene-understanding datasets (Cityscapes [8], Mapillary [32], BDD-100K [56], Dark Zurich [45]) fail to capture the full range of variations that exist in the world. Synthetic data is a popular alternative for generating diverse and bountiful views. GTAV [46], Synthia [43], and VisDA-C [38] are some of the widely-used synthetic datasets. These datasets can be curated using underlying simulators, such as GTAV [46] game engine and Carla [10] simulator and offer a cost-effective and scalable way to generate large amounts of labeled data under diverse conditions. Similar to SELMA [50] and SHIFT [49], we use Carla [10] as the underlying simulator for SkyScenes.
@@
Real-World Aerial Datasets. To support remote sensing applications, it is crucial to have access to datasets that offer aerial-specific views. Datasets such as GID [51], DeepGlobe [9], ISPRS2D [44], and FloodNet [40] primarily provide nadir perspectives and are designed for scene-recognition and understanding tasks. However, this study specifically focuses on lower altitudes, which are more relevant to UAVs, enabling object identification. Unfortunately, there is a scarcity of high-resolution real-world datasets based on UAV imagery emphasizing object identification. Existing urban scene datasets, like Aeroscapes [34], UAVid [27], VDD [1], UDD [6], UAVDT [11], VisDrone [57], Semantic Drones [20] and others, suffer from limited sizes and a lack of diverse images under different conditions. This limitation raises concerns regarding model robustness and generalization.
@@
Synthetic Aerial Datasets. Simulators can facilitate affordable, reliable, and quick collection of large synthetic aerial datasets, which aids in fast prototyping, improves real-world performance by enhancing robustness, and enables controlled studies on varied conditions. One such high-fidelity simulator, AirSim [48], used for development and testing of autonomous systems (in particular, aerial vehicles), is the foundation of several synthetic UAV-based datasets – MidAir [12], Espada [25], Tartan Air [53], UrbanScene3D [23] and VALID [5]. Carla [10] is another such open-source simulator that is the foundation of datasets like SynDrone[42]. However, these datasets fall short in capturing real-world irregularities, lack deterministic re-generation capabilities, controlled diversity in weather and daytime conditions, and exhibit skewed representation for certain classes (differences summarized in Table. 1). This restricts their ability to generalize well to real-world datasets and their usage as a diagnostic tool for studying the controlled effect of diversity on the performance of computer vision perception tasks. To enable such studies, SkyScenes offers images featuring varied scenes, diverse weather, daytime, altitude, and pitch variations while incorporating real-world irregularities and addressing skewed class representation along with simultaneous depth, semantic, and instance segmentation annotations.
@@ 3 SkyScenes
We curate SkyScenes using, Carla [10]333https://carla.org/ , which is a flexible and realistic open-source autonomous vehicle simulator. The simulator offers a wide range of sensors, environmental configurations, and varying rendering configurations. As noted earlier, we take several important considerations into account while curating SkyScenes images. These include strategies for obtaining diverse synthetic data and embedding real-world irregularities, avoiding correlated images, addressing skewed class representations, and more. In this section, we first discuss such desiderata and then describe our procedural image curation algorithm. Finally, we describe different aspects of the curated dataset.
@@
Before investigating the image curation pipeline, we first outline a set of desiderata taken into account while curating synthetic aerial images in SkyScenes.
@@
Viewpoint Reproducibility: Critical to understanding how models respond to changing conditions is the ability to evaluate them under scenarios where only one variable is altered. However, any effort to do so in the real-world would be uncontrolled, due to its dynamic (constantly changing) nature. In contrast, simulated data allows us to do so by providing control over image generation conditions. Unlike certain existing aerial datasets that do not support this feature (see Table. 1), we do so in SkyScenes by additionally storing comprehensive metadata for each viewpoint (and image), including details about camera world coordinates, orientation, and all movable/immovable actors and objects in the scene. We couple this with rigorous consistency checks for image generation that verify the number of actors, their location, sensor height, pitch, etc. This meticulous approach enables us to reproduce the same viewpoint under multiple conditions effortlessly.
2 .
Adequate Representation of Tail Classes: Unlike ground-view datasets, pixel distribution of classes in aerial images is substantially more long-tailed (see Fig. 2 (a); classes with smaller object size, humans). This substantial difference in class proportions severely affects the performance of tail classes in aerial datasets when compared to ground-view datasets (see Fig. 2 (b)), thus making visual recognition tasks harder. To counter this, we consider structured spawning of humans to ensure adequate representation (see Sec. 3.2).
@@
Adequate Height Variations: Aerial images are captured at different altitudes to meet specific needs. Lower altitudes (-m) are optimal for high-resolution photography and detailed inspections. Altitudes ranging from m-m strike a balance between fine-grained detail and a broader perspective, making them ideal for surveillance. Altitudes above m are suitable for capturing extensive areas, making them ideal for surveying and mapping. Existing datasets (synthetic or real) often focus on “specific” altitude ranges (see Table. 1, Image Capture columns), limiting their adaptability to different scenarios. With SkyScenes, our aim is to provide flexibility in altitude sampling, thus accommodating various real-world requirements. We curate SkyScenes images at heights of m, m, and m. Additionally, recognizing imperfections in real-world actuation, we induce slight jitter in the height values (m) to simulate realistic data sampling.
4.
Adequate Pitch Variations: Similar to height, aerial images can be captured from primary perspectives or pitch angles (): nadir (), oblique (), or forward () views (see Table. 1, Image Capture columns). The nadir view (directly perpendicular to the ground plane), preserves object scale while forward views are well-suited for tasks like UAV navigation and obstacle detection. Oblique views, on the other hand, capture objects from a side profile, aiding object recognition and providing valuable context and depth perspective often lost in nadir and forward views. To ensure widespread utility, SkyScenes data generation process is designed to support all these viewing angles, with a particular emphasis on oblique views (the most common one). Similar to height, pitch variations allow models trained on SkyScenes to generalize to different viewpoint variations. We use for oblique-views and introduce jitter () to mimic real-world data sampling.
5.
@@
6.
Adequate Weather & Daytime Variations: Training robust perception models using SkyScenes that generalize to unforeseen environmental conditions necessitates the curation of annotated images encompassing various weather and daytime scenarios. To accomplish this, we generate SkyScenes images from identical viewpoints under different variations – ClearNoon, ClearSunset, MidRainNoon, ClearNight, and CloudyNoon.444Note that Carla [10] provides such conditions but we use only such conditions in this preliminary version of SkyScenes. Generating images in different conditions from the same perspectives allows us to (1) leverage diverse data for improved generalization and (2) systematically investigate the susceptibility of trained models to variations in daytime and weather conditions.
7.